EVA Report #9 – May 13th

Crew 212 EVA Report 13-MAY-2019

EVA #09

Author: Vittorio Netti

Purpose of EVA:
Primary objective
Testing the first BVLOS mission of the "Palantir" fixes wing drone (Success)
Testing the flight attitude shifting capabilities of the "X-5" fixes wing VTOL drone (Partial Failure)

Measuring the Oxygen fluxus from the SpaceSuit under extensive stress (Partial Success)
Measuring the Oxygen level in blood after extensive stress in Spacesuit (Success)
Secondary objectives
Collecting aerial photos of the martian surface (Success)
Testing the assisting landing capabilities of the drone (Success)
Testing the ground control station during the auto mission mode (Success)

Start time: 16:00 AM

End time: 18:33 AM

Narrative:

The purpose of this EVA was to

The equipment was composed by:
– Palantir drone
– Ex Case 1 (pressurized)
– Ground Station Tripod

Since two experiments were foreseen for this EVA, two rovers were needed. Spirit presented anomalous battery consumption rates in the past days, so we decided to leave it at the Hab. Curiosity has always functioned nominally. Although Opportunity had overheating problems on last Sunday (Sol 0), it has been performing nominally in the past EVAs; while keeping the driving speed low, we decided to take it, as the EVA destination didn’t require a long drive (approx 15 minutes each way).

Opportunity departed from the hab with 100% of battery.
Curiosity departed from the hab with 100% of battery.

We departed from the Hab around 4:15 PM and drove towards north to free the mouse (captured last night) before Cowboy’s Corner; the speed was maintained low to avoid overheating problems with Oppi. After that, both rovers performed a U-turn and drove in the direction of URC North Site. At around 50 meters from the final objective, Opportunity stopped, and the battery overheating warning was on. Battery level at that point was 99%. We tried to start the engine many times, but the alarm was always on and didn’t let the engine start.

Since that we were on the experiment site, we decided to leave the rovers and start the experiment procedures, hoping that the Opportunity batteries can cool down during that time.

Together with Paolo we proceeded to assembly the drones and the GCS. At the same time, Zoe and Camilo proceeded to set a running track in order to perform their experiment under stress.

The "Palantir" drone has been hand-launched by Paolo and suddenly the control has been taken by the program, that started the automatic photosurvey of the area. During the mission "Palantir took more than 150photos, mapping more than 45 Hectare of terrain in 28 minutes.

Zoe and Camillo proceeded with the experiment, running through the track for 10 minutes with their suit on, on which they previously sealed the external openings.

After the landing, the Palantir has been recovered without damages. Together with Paolo we proceeded to assembly and arming the X-5 for his test.

The X-5 take-off happened at 17.45, but after the transition to the Plane mode the aerodynamic surfaces resulted not to be controllable enough from the software. We tried to perform a controlled soft landing without success. The plane suffered an impact to the ground that destroyed both the motor support plates.

At the end of experiments, since Opportunity was still off, we received authorization from the Director to use one rover to carry all four EVA crew members back to the Hab, leaving all the equipment in the overheated rover to have place for all the crewmember. The crashed drone has been left on his impact point in order to perform further analysis on the impact dynamic

Opportunity was retrieved later by the Director and an out-of-sim crew member. During the recovery, the X-5 has been photographed and then stored in the Opportunity trunk together with the rest of the equipment They were able to start the rover and drive to the Hab, where the rover stopped once again, luckily right in front of the Hab.


Destination:
URC North Site

Coordinates: (UTM NAD27 CONUS): 519012.00 E, 4251740.00 N, Zone 12S

EVA Participants: Vittorio Netti (Commander), Camillo Zorro (Medic), Zoe Townsend, Paolo Guardabasso

Road(s) and routes per MDRS Map: Cow Dung Road 0110

Mode of travel: driving

Operations Report – May 13th

Crew 212 Operations Report 13-MAY-2019

SOL: 8

Name of person filing report: Paolo Guardabasso

Non-nominal systems: Rover Opportunity, Upper deck water pump, Kitchen sink.

Notes on non-nominal systems: rover Opportunity failed during EVA (on the way to the second EVA objective). For more details, look at the notes about the rovers and the EVA report for EVA #9.

The upper deck pump is now activating for rather long times (tens of seconds) with all the sinks shut. The kitchen sink is leaking.

Generator (hours run last night): 11:15

On (last night): 7:45 PM

Off (this morning): 7:00 AM

Solar— SOC 61% at 7:45 PM

Diesel Reading – 88 %

Propane Reading – 46%

Ethanol Free Gasoline – 4 gallons

Water (Axillary tank) – 0 gallons

Water (static tank) – 380 gallons

Axillary to Static tank transfer – no

Gallons transferred: ~200

Water in GreenHab – 137 gallons

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 143583.1

Toilet tank emptied: No

Deimos rover used: no (out of order)

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: No

Hours: 92.8h

Beginning charge:

Ending charge:

Currently charging: Yes

Opportunity rover used: Yes

Hours: 59.5h

Beginning charge: 100

Ending charge: 98

Currently charging: Yes

Curiosity rover used: Yes

Hours: 95.6h

Beginning charge: 100%

Ending charge: 97%

Currently charging: Yes

Notes on rovers: Opportunity rover failed during EVA 9 because of overheating of battery (more details in EVA report); it was not possible to restart the engine until . Problem might be cause by lack of distilled water in radiator. Spirit rover battery has had anomalous consumption rates in the past days. A major rover check will be run tomorrow during an Engineering EVA.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3)

Reason for use:

Oil Added?

ATV Fuel Used: Gals

# Hours the ATVs were Used today:

Notes on ATVs:

HabCar used and why, where? The HabCar was used to reach the location were Opportunity was left after it failed during EVA 9.

CrewCar used and why, where? Off Campus

General notes and comments: We captured another mouse last night, and it was set free today during EVA 9 at Cowboy’s Corner. EVA 9 presented several challenges (see EVA report).

Summary of internet: Nothing to report.

Summary of suits and radios: some suits batteries are at end of life, they lose charge very quickly (see spacesuit engineer report).

Summary of Hab operations: The ceiling in the commander’s room is compromised by water and might have a localized collapse if weight is applied to it, a bucket had to be placed close to the bed. A mouse was found in the lower deck. Leaking of the sink in the kitchen has been found, and this could be the cause of the leaking reported in a previous report.

Summary of GreenHab operations: Nothing to report.

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: what is the level of usage of the potable water filter? It is quite brown at the moment (see photo attached). Should it be replaced together with the pump tomorrow night?

Spacesuit Report – May 13th

Crew 212-LATAM-III

Submitted by Spacesuit Engineer: Camilo Andrés Zorro Mendoza

Spacesuit report

Sol: 8

Sim suit used EVA: 1, 8, 3 and 5.
used for 2 and 1/2 hour

Sim suit: 1
Voltage level before EVA: 12.8 V
Voltage level after EVA:
12.7 V

Sim suit: 8
Voltage level before EVA: 12.8 V
Voltage level after EVA:
4.1 V

Sim suit: 3
Voltage level before EVA:
13 V
Voltage level after EVA:
12.8 V

Sim suit: 5
Voltage level before EVA: 13.2 V
Voltage level after EVA:
7.6 V

General commentaries

It is need get 10 new batteries for the radios. It was getting fake charge. 4 radios with good battery

Journalist Report – May 13th

Today was Sol 8.

Today started with a relaxed morning, the crew members worked on their own projects. There was the continuation of work on processing the photographs, more progress was made to recover the software of the rover with another 2am conference with the UK, and another batch of bacteria was mixed with soil to re-inoculate the seeds, the previous equipment has been sterilised and the bacteria has been killed.

An EVA was taken today which was 2 ½ hours long, it was an eventful EVA that was mixed with an engineering EVA. Firstly, we delivered another Mars mouse to the release point, after we went to the experiment site and started the two experiments. The first was to test the oxygen levels in the suits; a circuit was completed to show the differences in the oxygen levels. However, due to the air holes and lack of sealing a new experiment must be proposed. The drone experiment continued, with the testing the X-5 and the Palantir. A failure occurred in the Palantir during piloting, the damage seems unrepairable. Unfortunately, the EVA finished with a failure of Opportunity. Opportunity had to be towed back to the Hab during and engineering EVA.

Sol Summary – May 13th

Crew 212 Sol 8 Summary Report 13-MAY-2019

Sol: 8

Summary Title: Sun is really strong right now

Author’s name: Marlen Castillo Vilcahuaman

Mission Status: A-ok, one rover is down

Sol Activity Summary: For today, we had an EVA in the afternoon. Opportunity has overheated and had to be left on the road. Right now, one crew member and the director are going to retrieve it.

Look Ahead Plan: Right now, our priority is to a) discuss the issues in this EVA and b) retrieve the rover. We’ll probably be doing this.

Anomalies in work: Opportunity overheated and the crew members in EVA had to return without it. We’re also having trouble with the pump, because it has been leaking more water than usual.

Weather: Sun is strong.

Crew Physical Status: Although Opportunity overheated, crew members had water with them so they arrived from the EVA safe and hydrated.

EVA: 1

Reports to be file: Operations Report, GreenHab Report, Journalist Report, Daily Photo Report, EVA report, Astronomy report (please, wait while we’re trying to solve the trouble with Opportunity)

Support Requested: None.

Greenhab Report – May 13th

Crew 212 GreenHab Report – 13-May-2019: SOL 8

GreenHab Officer: Hector Palomeque

Environmental control: Ambient with the door open + fan turned up to 3 (max) from 12:00 to 19:00

Shade cloth 80%: on

Average temperatures:
Low: 20 °C
High: 43 °C

Hours of supplemental light: 0

Daily water usage for crops: 7 gallons

Water in Blue Tank: 137 gallons

Time(s) of watering for crops: 12:00

Changes to crops: All plants look healthy. All tomato plants continue blooming.

Narrative: The GreenHab fan had to be turned on because of the indoor temperature has reached +40 °C at noon. Plants were watered once at noon.

Basil harvested to add some flavor to the Crew’s lunch.

Harvested: Basil 20g

Support/supplies needed: None.

Astronomy Report – May 13th

Astronomy Report
Name: Mariona Badenas
Crew: 212
Date: 13 May 2019

MDRS ROBOTIC OBSERVATORY
Robotic Telescope Requested (choose one): MDRS-WF

** Object Viewed #1: IC5067 (Pelican Nebula) –> Observations completed.

Problems Encountered: N/A.

** Object Viewed #2: Veil Nebula

Problems Encountered: N/A –> Observations will continue tonight provided the weather is good.

** Object Viewed #3: Iris Nebula
Problems Encountered: N/A –> Observations will continue tonight provided the weather is good.

** (More objects to be Viewed): NGC6992, IC1396
Problems Encountered: N/A –> Observations of these objects have not begun yet.

MUSK OBSERVATORY
** Solar Features Observed: Observed the sun through the eyepiece, but couldn’t identify features on the surface due to problems with the camera settings and the ICap software.

Problems Encountered: The camera driver used for ICap 2.3 was not correct. This problem has already been solved. I will try to take pictures of the solar corona tomorrow, following Peter’s new instructions (private communication).